GETTING TO THE CORE OF WATER

\underline{DAVID W. SCHWENKE}, MS T27B-1, Moffett Field, CA 94035-1000.

In the famous water potential of Partridge and Schwenke (PS) [J.Chem.Phys 106 4618 (1997)], the critical ingredient for the very successful fit to experimental data was the inclusion of core-valence correlation. Other corrections to the potential, such as scalar relativity and Born-Oppenheimer breakdown, are of non-negligible size,[Polyansky et al. Science 299 539 (2003)] but can be easily modeled empirically. A drawback of the PS calculation of the core-valence correction was it was carried out using just one basis set, and that the quality of the basis set was unknown. We have developed a new method for judging the quality of core-valence basis sets, and this has enabled us to optimize a new hierarchy of basis sets for treating core-valence correlation in a much more systematic manner. These new basis sets will be used to provide a new, more accurate, calculation of the core-valence correction to the potential, dipole moment and spectrum of water.